# Faster and Better 3D Splatting via Group Training

### v1

@@ 4.4 Discussion
$\bullet$ Why faser? Group Training significantly accelerates reconstruction primarily by reducing the number of Under-training Gaussian primitives, thereby decreasing the rendering burden on the model. This speedup is observed even when the final model size is larger than the baseline. In alignment with our earlier analysis in Sec. 3.2, the OPS strategy also contributes notably to this speedup effect. During the Gaussian densification phase, Group Training with OPS preserves high-opacity Gaussians within the Under-Training Group, without which could lead to instances of both under-reconstruction and over-reconstruction [12, 29]. These issues can lead to intense Gaussian densification, as experimentally analyzed in Sec. 3.2. As a result, models reconstructed with Group Training and OPS contain fewer redundant Gaussians, leading to smaller model sizes and faster reconstruction overall. Additionally, OPS prioritizes sampling high-opacity Gaussians, which rapidly drives opacity saturation, which also discussed in Sec. 3.2. Focusing on these most impactful Gaussians reduces the number of Gaussians needed for splatting, further enhancing training and rendering efficiency.
@@ 5 Conclusion
Appendix A Efficiency And Effectiveness For Various Sampling Strategies
$p_{i}=\frac{\theta_{i}}{\sum_{i=1}^{N}\theta_{i}},\vspace{-2pt}$ (5)
$\theta_{i}=\alpha_{i}\cdot V_{i},$ (6)
@@ Appendix A Efficiency And Effectiveness For Various Sampling Strategies
Appendix B Detailed Experimental Results for All Scenes
We present the reconstruction results for all scenes using Group Training with Random Sampling (RS) and Opacity-based Prioritized Sampling (OPS), evaluated on 3D Gaussian Splatting (3DGS) [12] and Mip-Splatting [30]. The detailed results are provided in  Tabs. 11, 9, 8, 10, 12, 13, 15 and 14.

### v2

@@ 4.5 Discussion
$\bullet$ Why faser? Group Training significantly accelerates reconstruction primarily by reducing the number of Under-training Gaussian primitives, thereby decreasing the rendering burden on the model. This speedup is observed even when the final model size is larger than the baseline. In alignment with our earlier analysis in Sec. 3.2, the OPS also contributes notably to this speedup effect. During the densification phase, Group Training with OPS preserves high-opacity Gaussians within the Under-Training Group, without which could lead to instances of both under-reconstruction and over-reconstruction [12, 29]. These issues can lead to redundant densification. As a result, models reconstructed with Group Training and OPS contain fewer redundant Gaussians, leading to smaller model sizes and faster reconstruction overall. Additionally, OPS prioritizes sampling high-opacity Gaussians, which rapidly drives $\alpha$ saturation, which also discussed in Sec. 3.2. Focusing on these most impactful Gaussians reduces the number of Gaussians needed for splatting, further enhancing training and rendering efficiency.
@@ 5 Conclusion
Appendix A Proof of Property 1: Opacity-based Effective Gaussians Densification
Under the assumptions of mutual independence between Gaussian attributes and intra-primitive parameter independence, the partial derivatives $[\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\frac{\partial\Delta y}{\partial y_{\mathrm{m}}}]$ for any arbitrary Gaussian admit the following derivation:
[…]
$\frac{\partial L}{\partial x_{\mathrm{m}}}&=\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial G^{\mathrm{2D}}_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}}\\ &=\sum\limits_{\mathrm{pixel}}o_{m}\frac{\partial L}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}}\\ &=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\end{aligned}$
$\frac{\partial L}{\partial y_{\mathrm{m}}}&=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}\frac{\partial\Delta y}{\partial y_{\mathrm{m}}},\end{aligned}$
where $[\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\frac{\partial\Delta y}{\partial y_{\mathrm{m}}}]$ remain constant parameters determined by the resolution $[W,H]$; $[\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x},\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}]$ derive from the scale, the rotation and the world coordinates of Gaussian primitives (independent of their opacity); and $\frac{\partial L}{\partial\hat{C}}$ represents the loss gradient with respect to the current pixel value.
Given that $\hat{C}$ is formulated as the composite rendering of $N$ Gaussians in Eq. 14, the derivative $\frac{\partial\hat{C}}{\partial\alpha_{m}}$ admits computation via Eq. 15.
$\hat{C}&=\underbrace{\sum\limits_{i=1}^{m-1}\alpha_{i}c_{i}\prod\limits_{j=1}^{i-1}{(1-\alpha_{j})}}_{\mathrm{Before\,Gaussian}\,m}+\underbrace{\alpha_{m}c_{m}\prod\limits_{j=1}^{m-1}{(1-\alpha_{j})}}_{\mathrm{Gaussian}\,m}\\ &+\underbrace{\sum\limits_{i=m+1}^{N}\overbrace{\prod\limits_{j=1}^{m-1}{(1-\alpha_{j})}}^{\mathrm{Before\,G\,}m}\alpha_{i}c_{i}(1-\alpha_{m})\overbrace{\prod\limits_{j=m+1}^{i-1}{(1-\alpha_{j})}}^{\mathrm{After\,G\,}m}}_{\mathrm{After\,Gaussian\,}m}\\ &+\underbrace{\prod\limits_{i=1}^{N}(1-\alpha_{i})c_{\mathrm{bg}}}_{\mathrm{background}},\end{aligned}$
$\frac{\partial\hat{C}}{\partial\alpha_{m}}=\overbrace{\prod\limits_{j=1}^{m-1}{(1-\alpha_{j})}}^{\mathrm{Before\,G\,}m}\left[c_{m}-\sum\limits_{i=m+1}^{N}\alpha_{i}c_{i}\overbrace{\prod\limits_{j=m+1}^{i-1}{(1-\alpha_{j})}}^{\mathrm{After\,G\,}m}\right]-\frac{c_{\mathrm{bg}}T_{N}}{1-\alpha_{m}}\end{aligned}$
Subsequently, the mathematical expectation of this derivative is formally established through Eq. 16
$\mathbb{E}\left[\frac{\partial\hat{C}}{\partial\alpha_{m}}\right]&=\overbrace{(1-\alpha_{0})^{m-1}}^{\mathrm{Before\,G\,}m}\left\{c_{0}-c_{0}\alpha_{0}\sum\limits_{i=m+1}^{N}\mathbb{E}\left[\overbrace{\prod\limits_{j=m+1}^{i-1}{(1-\alpha_{j})}}^{\mathrm{After\,G\,}m}\right]\right\}-\frac{c_{\mathrm{bg}}T_{\mathrm{sta.}}}{1-\alpha_{0}}\\ &=(1-\alpha_{0})^{m-1}\left[c_{0}-c_{0}\alpha_{0}\sum\limits_{i=m+1}^{N}(1-\alpha_{0})^{i-m-1}\right]-\frac{c_{\mathrm{bg}}T_{\mathrm{saturation}}}{1-\alpha_{0}}\\ &=\frac{(c_{0}-c_{\mathrm{bg}})T_{\mathrm{saturation}}}{1-\alpha_{0}}\\ &=\frac{(c_{0}-c_{\mathrm{bg}})T_{\mathrm{saturation}}}{1-\mathbb{E}\left[o_{i}\right]\mathbb{E}\left[G_{i}\right]},\end{aligned}$
@@ Appendix A Proof of Property 1: Opacity-based Effective Gaussians Densification
Mip-NeRF360 [1] Tanks&Temples [13] Deep Blending [9] Blender [20]
PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$ PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$ PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$ PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$
3D-GS  [12] 27.45 26.7 23.70 15.0 29.59 23.9 33.77 6.1
+Group Training 27.56 19.6 27% 23.85 11.0 27% 29.75 16.9 29% 33.81 4.8 21%
Mini-Splatting  [6] 27.27 20.7 23.26 12.6 29.95 17.8 31.60 10.0
+Group Training 27.25 17.9 13% 23.10 9.9 21% 29.85 14.7 17% 31.98 8.4 16%
LightGaussian  [5] 27.06 27.5 23.09 16.1 27.28 25.9 32.95 6.1
+Group Training 27.34 20.5 25% 23.55 11.9 26% 28.50 19.0 27% 33.18 4.6 24%
Table 8: Quantitative comparisons on different baselines and datasets. Group Training with 3DGS achieves faster reconstruction speed than Mini-Splatting across all datasets. Furthermore, Group Training demonstrates consistent acceleration effects on both 3DGS acceleration model (13$\sim$21speedup on Mini-Splatting [6]) and compression model (24$\sim$27speedup on LightGaussian [5]).Accel. = Acceleration Ratio in training time compared to the baseline.
Appendix B Efficiency And Effectiveness For Various Sampling Strategies
$p_{i}=\frac{\theta_{i}}{\sum_{i=1}^{N}\theta_{i}},\vspace{-2pt}$ (17)
$\theta_{i}=\alpha_{i}\cdot V_{i},$ (18)
Appendix C Methodological Applicability
We perform comparative validation across two distinct 3DGS architectures: an acceleration-optimized model [6] and a compression-focused LightGaussian [5].
Empirical results demonstrate Group-Training’s consistent efficacy across dataset scales, particularly evidenced by reduced temporal overhead in the Blender [20], as shown in Tab. 8. Crucially, our method synergistically integrates with existing acceleration techniques like Mini-Splatting [6], achieving compounded acceleration gains while providing sustained acceleration for compressed models with concurrent fidelity enhancement.
Appendix D Temporal Evolution of Under-Training Gaussian Primitives.
Figure 8: Comparsion of Under-Training Gaussian Primitives. Our Group-Training methodology selectively trains a subset of Gaussian primitives, demonstrating enhanced computational efficiency while mitigating loss of potentially critical points during opacity reset operations.
We visually compare the quantitative differences in under-training Gaussian primitives between Group-Training and 3DGS during scene reconstruction in Fig. 8. 3DGS with Group-Training reduces the training overhead by avoiding full optimization of all Gaussian primitives. Furthermore, during each opacity reset operation, the proposed method retains a higher proportion of geometrically significant primitives compared to baseline. These retained elements, despite their low-opacity values, preserve critical structural information that contributes to scene geometry fidelity.
Appendix E Detailed Experimental Results for All Scenes
We present the reconstruction results for all scenes using Group Training with Random Sampling (RS) and Opacity-based Prioritized Sampling (OPS), evaluated on 3D Gaussian Splatting (3DGS) [12] and Mip-Splatting [30]. The detailed results are provided in  Tabs. 12, 10, 9, 11, 13, 14, 16 and 15.